# Harry Yang1, Linfeng Zhang3, Qifeng Chen1🖂, Yue Ma1🖂

### v2

@@ 8 Limitations and Future Work
8.1 Robustness to Prompt and Hyperparameters
First, the framework exhibits high sensitivity to prompt description. Since the model’s editing behavior is driven entirely by prompt-guided inversion and denoising trajectories, slight variations in wording or structure can lead to noticeably different editing outcomes. This makes it difficult to guarantee consistent performance across prompts, especially in open-ended or ambiguous settings.
Second, as a training-free method, our approach relies on carefully tuned hyperparameters (e.g., injection timing, control strength). However, no single configuration generalizes optimally across all tasks and image categories. The absence of a learnable mechanism makes it challenging to adaptively adjust parameters for different editing scenarios, which may limit flexibility in real-world applications. However, in practical applications, the robustness of the framework is required to be concerned [25]. Future work could explore developing prompt-robust techniques that maintain semantic fidelity across a broader range of user inputs.
@@ 8.2 Extending to Video Editing
We also explore extending our shape-aware editing framework to the video domain using Wan 2.1 [48], an open-source video generation model utilized Rectified Flow. While our method can in principle be applied to all frames, we find that the temporal dimension introduces a major challenge, where the TDM becomes much less stable and effective when extended across time, as shown in Figure 8. In particular, the spatial editing regions indicated by TDM often fluctuate across frames, leading to inconsistent or incomplete transformations in the resulting video. Since a well-defined and temporally consistent TDM is crucial for successful editing, future work may consider strategies such as temporally-aware TDM construction, or explicit disentanglement of spatial and temporal components in the denoising trajectory.
@@ 9 More Editing Results
We present additional editing results in Figure 10, all conducted on images in PIEBench and ReShapeBench.

### v4

@@ 10 Limitations and Future Work
10.1 Failure Cases
Our method can be sensitive to prompt ambiguity and imprecise editing instructions. Since the editing behavior is driven entirely by prompt-guided inversion and denoising trajectories, the quality of the editing outcome depends on how clearly the intended modification is specified in the text prompt. When editing instructions are vague, lack clear semantic targets, or have low discriminative specificity, the model may struggle to determine where and how strongly to apply the modification. This often leads to weak, diffuse, or inconsistent edits, particularly in cases where the intended change is not explicitly specified by the prompt. For example, prompts that describe abstract transformations or rely on implicit assumptions about the editing target may result in edits that do not match user expectations (see Figure 11). Clear and well-defined editing descriptions that explicitly identify the object to be modified and the desired transformation are therefore important for reliable performance.
Figure 11: Failure Case
@@ 10.2 Extending to Video Editing
Video generation [40, 45, 46, 43, 38, 42, 44, 11, 78, 25, 16, 67, 65, 53, 56, 89, 88, 64, 55, 57, 54, 66, 85, 83, 82, 81, 84, 14, 5] is an important research direction for practical application. We also explore extending our shape-aware editing framework to the video domain using Wan 2.1 [71], an open-source video generation model that uses Rectified Flow. While our method can in principle be applied to all frames, we find that the temporal dimension introduces a major challenge, where the TDM becomes much less stable and effective when extended across time, as shown in Figure 12. In particular, the spatial editing regions indicated by TDM often fluctuate across frames, leading to inconsistent or incomplete transformations in the resulting video. Since a well-defined and temporally consistent TDM is crucial for successful editing, future work may consider strategies such as temporally-aware TDM construction, or explicit disentanglement of spatial and temporal components in the denoising trajectory.
10.3 Extending to Agent-based Editing
An important future direction is to extend our framework toward an agent-based video editing paradigm [26, 76, 75, 74, 27, 28, 31, 91, 92, 93, 8, 7]. Instead of treating editing as a single-pass generation problem, we envision a multi-agent system in which specialised agents are responsible for temporal consistency verification, semantic alignment, motion-aware refinement, and user-intent interpretation. Through iterative reasoning and tool invocation (e.g., tracking, depth estimation, or diffusion-based refinement modules), such agents could decompose complex editing instructions into structured sub-tasks and dynamically adjust intermediate results. This formulation would enable adaptive long-horizon editing, reduce error accumulation across frames, and improve controllability under ambiguous prompts. Moreover, integrating memory mechanisms for cross-frame state tracking may further enhance temporal coherence and enable interactive, feedback-driven video manipulation.
@@ 11 More Editing Results
We present additional shape-aware editing results in Figure 15 and Figure 16. We also present general task editing results in Figure 17.
12 More Editing Results
We present additional shape-aware editing results in Figure 15 and Figure 16. We also present general task editing results in Figure 17.
Figure 13: Single-Object Cases in ReShapeBench
Figure 14: Multi-Object Cases in ReShapeBench
Image Prompts
Source Prompt: A single Air Jordan 1 high-top in University Blue rests elegantly atop a weathered tree stump, its bold panels and crisp lines catching the golden afternoon light. The rich blue leather, white toe box, black laces, and dark grey Swoosh stand out sharply against the rough bark beneath, every stitch and texture rendered with photorealistic clarity. Behind it, a soft blur of ivy leaves and tangled branches melts into warm bokeh, hinting at a quiet forest edge bathed in dappled sunlight. The entire scene feels effortlessly stylish — where streetwear meets nature, serene yet striking, captured with cinematic depth and organic warmth.
Edit Prompt 1: A sleek tabby cat rests elegantly atop a weathered tree stump, its graceful curves and soft fur catching the golden afternoon light. The rich orange-brown coat, white chest patch, black-tipped ears, and whiskers stand out sharply against the rough bark beneath, every strand rendered with photorealistic clarity. Behind it, a soft blur of ivy leaves and tangled branches melts into warm bokeh, hinting at a quiet forest edge bathed in dappled sunlight. The entire scene feels gently alive — where wild nature cradles quiet companionship, serene and soulful, captured with cinematic depth and organic warmth.
Edit Prompt 2: A delicate porcelain teacup with matching saucer rests gracefully atop a weathered tree stump, its fine floral patterns and golden rim catching the golden afternoon light. The white ceramic surface, intricate pink blossoms, green leaves, and gilded edges stand out sharply against the rough bark beneath, every detail rendered with photorealistic clarity. Behind it, a soft blur of ivy leaves and tangled branches melts into warm bokeh, hinting at a quiet garden edge bathed in dappled sunlight. The entire scene feels timelessly elegant — where refined craftsmanship meets rustic nature, serene yet sophisticated, captured with cinematic depth and organic warmth.
Table 5: Example 1: Image–prompt pairs in ReShapeBench
Image Prompts
Source Prompt: A steaming cup of latte with delicate leaf art sits beside two slices of avocado toast topped with runny yolks and fresh herbs, all bathed in soft morning light from a nearby window. The creamy foam, golden crust, vibrant green leaves, and glistening egg yolk stand out sharply against the rustic ceramic plate, every crumb and droplet rendered with photorealistic clarity. Behind it, a gentle blur of lush green foliage outside the glass pane melts into warm bokeh, hinting at a quiet garden waking under diffused daylight. The entire scene feels gently inviting — where simple pleasures meet slow mornings, serene yet richly textured, captured with cinematic depth and organic warmth.
Edit Prompt 1: A delicate sprig of white wildflowers rests gently beside two slices of avocado toast topped with runny yolks and fresh herbs, all bathed in soft morning light from a nearby window. The fragile petals, dewy leaves, golden crust, and glistening egg yolk stand out sharply against the rustic ceramic plate, every texture rendered with photorealistic clarity. Behind it, a gentle blur of lush green foliage outside the glass pane melts into warm bokeh, hinting at a quiet garden waking under diffused daylight. The entire scene feels tenderly alive — where nature’s grace meets simple nourishment, serene yet richly textured, captured with cinematic depth and organic warmth.
Edit Prompt 2: A ripe yellow banana rests casually beside two slices of avocado toast topped with runny yolks and fresh herbs, all bathed in soft morning light from a nearby window. The smooth peel, gentle curve, golden hue, and glistening egg yolk stand out sharply against the rustic ceramic plate, every texture rendered with photorealistic clarity. Behind it, a gentle blur of lush green foliage outside the glass pane melts into warm bokeh, hinting at a quiet garden waking under diffused daylight. The entire scene feels playfully alive — where everyday fruit meets wholesome nourishment, serene yet subtly whimsical, captured with cinematic depth and organic warmth.
Table 6: Example 2: Image–prompt pairs in ReShapeBench
Image Prompts
Source Prompt: A young boy crouches low on a quiet street, one hand resting gently on a worn soccer ball, his gaze steady and thoughtful under the soft glow of late afternoon light. His teal striped sweater, red shorts with white stripes, scuffed sneakers, and the cracked leather surface of the ball stand out sharply against the rough asphalt beneath, every thread, crease, and stitch rendered with photorealistic clarity. Behind him, a gentle blur of weathered brick walls and distant buildings melts into warm bokeh, hinting at a humble neighborhood bathed in golden-hour haze. The entire scene feels quietly powerful — where childhood dreams meet everyday resilience, serene yet deeply human, captured with cinematic depth and organic warmth.
Edit Prompt 1: A young boy crouches low on a quiet street, one hand resting gently on a worn leather backpack beside his knee, his gaze steady and thoughtful under the soft glow of late afternoon light. His teal striped sweater, red shorts with white stripes, scuffed sneakers, and the frayed straps and faded stitching of the backpack all stand out sharply against the rough asphalt beneath, every texture rendered with photorealistic clarity. Behind him, a gentle blur of weathered brick walls and distant buildings melts into warm bokeh, hinting at a humble neighborhood bathed in golden-hour haze. The entire scene feels quietly nostalgic — where school days meet quiet contemplation, serene yet deeply human, captured with cinematic depth and organic warmth.
Edit Prompt 2: A young boy crouches low on a quiet street, one hand resting gently on the back of a sleepy dog curled beside his sneaker, his gaze steady and thoughtful under the soft glow of late afternoon light. His teal striped sweater, red shorts with white stripes, scuffed sneakers, and the soft fur, floppy ears, and relaxed posture of the dog all stand out sharply against the rough asphalt beneath, every texture rendered with photorealistic clarity. Behind him, a gentle blur of weathered brick walls and distant buildings melts into warm bokeh, hinting at a humble neighborhood bathed in golden-hour haze. The entire scene feels tenderly alive — where childhood quietness meets loyal companionship, serene yet deeply human, captured with cinematic depth and organic warmth.
Table 7: Example 3: Image–prompt pairs in ReShapeBench
Figure 15: Additional Editing Results
Figure 16: Additional Editing Results
Figure 17: Additional Editing Results